As an airplane descends toward an airport, it drops a vertical distance of 23 m and moves forward a horizontal distance of 330 m .What is the distance covered by the plane during this time?
1
Expert's answer
2018-09-17T15:08:09-0400
v^2=〖24〗^2+〖10〗^2
v=√(〖24〗^2+〖10〗^2 )=√(2^2⋅〖12〗^2+2^2⋅5^2 )=2⋅√(〖12〗^2+5^2 )=2⋅√(144+25)=2⋅√169=2⋅13=26 m⁄s